Your portfolio has a beta of 1.23. The portfolio consists of 15 percent U.S. Treasury bills, 30 percent in stock A, and 55 percent in stock B. Stock A has a risk-level equivalent to that of the overall market. What is the beta of stock B?

1.69

1.10

1.40

1.29

0.55
Answer #1

Portfolio beta=Respective beta*Respective weight

1.23=(0.15*0)+(0.3*1)+(0.55*Beta of stock B)[Beta of market=1;Beta of risk-free assets=0]

1.23=0.3+(0.55*Beta of stock B)

Beta of stock B=(1.23-0.3)/0.55

=1.69(Approx).